# LeTac-MPC: Learning Model Predictive Control for Tactile-reactive Grasping

### v1

@@ III-C Training
For this MPC layer, we choose $\mathbf{A}_{f}\in\mathbb{R}^{M}$ and $\mathbf{Q}_{f}\in\mathbb{R}^{M\times M}$ as the parameters we want to learn, since they cannot be derived from the model-based formulation. On the other hand, the scalars $Q_{v}$, $Q_{a}$, and $P$ are model-based control parameters that are directly related to the response and convergence speed of the MPC. Therefore, we do not consider them as parameters to learn and can assign reasonable values directly. Additionally, it is important to note that $Q_{v}$, $Q_{a}$, and $P$ must be positive.
Optimization problem (III-B) is a quadratic program (QP). The work in [42] demonstrates that a QP can be represented as a differentiable layer. In addition, a method is proposed for solving multiple QPs in batch form in [42]. Therefore, the NN shown in Fig. 1 can perform forward pass and backpropagation in batch form if the optimization problem (III-B) is feasible throughout the training process. To ensure feasibility, the resulting QP’s Hessian matrix must remain symmetric positive definite when $\mathbf{A}_{f}$ and $\mathbf{Q}_{f}$ change during training [43].
Regarding the feasibility of our proposed differentiable MPC layer, we state the following theorem:
@@
Since the optimization problem (III-B) has no constraints besides (22), the resulting QP is unconstrained. Therefore, to prove the theorem, it is sufficient to demonstrate that the Hessian matrix of the resulting QP is symmetric positive definite [43]. The resulting Hessian matrix from (III-B) is
@@
Because $Q_{v},Q_{a},P>0$ and $\mathbf{Q}_{f}$ is symmetric positive definite, $\bar{\mathbf{Q}}$ and $\bar{\mathbf{Q}}_{a}$ are symmetric positive definite. Therefore, by equation (25), we can see that the Hessian matrix is symmetric positive semi-definite.
In order to prove Hessian matrix is symmetric positive definite, we can prove rank($\bar{\mathbf{S}}$$=N$. Because $\mathbf{B}\neq\mathbf{0}$, we can clearly see from equation (30) that rank($\bar{\mathbf{S}}$$=N$. Therefore, we have shown that the Hessian matrix of the resulting QP is symmetric positive definite, which completes the proof of the theorem. ∎
Theorem 1 demonstrates that if we ensure $\mathbf{Q}_{f}$ is symmetric positive definite, the selection of $N$ and $M$, and uncertainty in $\mathbf{f}_{n}$ and $\mathbf{A}_{f}$ will not affect the QP’s feasibility. As a result, we can freely tune the parameters and train the model stably. To ensure that $\mathbf{Q}_{f}$ remains symmetric positive definite during training, we utilize a Cholesky factorization
@@
Although Theorem 1 shows that the selection of prediction horizon $N$ and embedding dimension $M$ does not affect the feasibility of the QP, it is still important to choose appropriate values for these parameters. A very small $N$ and $M$ can negatively impact the network’s ability to extract useful information and generalize well. Conversely, setting $N$ and $M$ to be too large can slow down the optimization problem’s solution due to the increase in dimensionality when implementing the trained model as a controller, which can result in a low control frequency. Moreover, when $N$ is excessively large, the linear MPC layer may result in a large prediction error. Therefore, the choice of $N$ and $M$ should balance the expressiveness of the model and the controller’s computational efficiency. In our implementation, we choose $N=15$ and $M=20$ to strike this balance.
@@
Figure 2: For each type of material, bottom left is the tactile raw image overlapping with the marker tracking when grasping the corresponding block and bottom right is the tactile raw image without marker tracking. These tactile images are collected under the same grasping force. Our goal here is to show that different materials have different physical properties, which leads to different features in the tactile image under the same grasping conditions. Therefore, we do not estimate the specific values of tangential force here.
@@
Figure 3: Left: data collection setup. Right: automated data collection pipeline.
@@ III-D Data Collection
We selected four different types of materials for data collection, as illustrated in Fig. 2. It is important to mention that we chose only standardized blocks without surface texture, and do not require specific sizes, as long as they fit the gripper. By doing so, the data collection process is simplified. we show in Section V that our proposed network model can learn from these standardized blocks and generalize to other daily objects with different textures, sizes, shapes, and physical properties.
@@
These blocks are used as grasped objects to capture tactile raw images. The tactile feedback of these blocks varies depending on their unique physical properties. As shown in Fig. 2, we can see the differences in tactile feedback between these 4 blocks.
@@
The contact area of hard materials is smaller but more distinct, while the contact area of soft materials is larger but subtler. This is because of the difference of the objects’ stiffness. When objects with different stiffness are grasped under the same applied force, the deformation of the grasped objects and the resulting deformation on the GelSight surface will be different, which results in different tactile raw images.
@@
The 3D printing block serves the role of a rigid object. We utilized a Markforged Onyx printer, which prints using micro carbon fiber filled nylon. This material is Onyx composite material. Its tensile modulus is 2.4 Gpa, and its flexural modulus is 3 Gpa [44]. When printed with a block form factor, it behaves rigid when grasped by our grippers. Compared with the hard block made by 3D printing, hard rubber is softer, so the contact area is larger. Additionally, the contact area of gel is the most subtle. The gel is made of XP-565 with shore A hardness of 27 [45], which is approximate 0.0009942 GPa of the Young’s modulus [46]. The stiffness of the softer rubber is between the gel and hard rubber. Finally, the difference in marker displacements can be attributed to the fact that materials with higher surface friction coefficients tend to induce more pronounced displacement on the GelSight’s surface.
@@
We believe that when other people selecting materials to construct a dataset, following our approach of choosing one hard material (such as rigid 3D printed objects or wood) and then selecting materials with progressively lower stiffness, up to a total of four materials, would be an effective process for dataset construction.
@@
Our network input only consists of raw images. Since marker displacements are actually obtained by optical flow from raw images, the raw images themselves already contain this information. To collect data for training the NN shown in Fig. 1, we propose an automatic data collection pipeline utilizing a dual-arm setup. We use Franka Panda robotic arm. One of the arms serves as the leader arm, and we control it by make it follow specific trajectories. The other arm serves as the follower arm, and we implement an impedance controller for it. The WSG 50-110 grippers are attached to the end-effectors of both arms, with one finger on the follower arm mounting the GelSight tactile sensor.
@@
Next, the follower gripper is relaxed gradually with a velocity of 4.5 mm/s, and we record the width of the follower gripper $p_{n}$ as well as the raw image of the tactile feedback at 60 Hz. When the follower gripper’s grasping force applied to the object cannot provide a maximum static friction force greater than the external force from the impedance controller, slipping occurs between the follower gripper and the object. This results in the follower arm moving back to the equilibrium point of the impedance controller. We record the width of the follower gripper when slipping occurs as the label for all raw images recorded during the entire trial, denoted as $p_{\text{slip}}$.
Ideally, we hope the leader gripper maintains an absolutely tight grasping throughout the process and there is no slipping between the leader gripper and the grasped object. To achieve that, we maximize the grasping force of the leader gripper on the data collection block and enhance the friction between the leader gripper’s fingers and the block by adding a gel layer on the finger. Naturally, we could not guarantee that there would be no movement between the object and the leader gripper during extended periods of data collection. However, our setup ensures that any movement, if it occurred, is minimal and does not affect the data collection. Therefore, when the leader gripper stays static, the grasped object stays static and we set the velocity of the end-effector of the follower arm to be greater than 5 mm/s as the condition for indicating slipping. Once slipping occurs for 0.2 s, we stop recording the raw images. The reasons of selecting 0.2 s are as follows. For each material, as shown in Fig. 4, when the external force varies from 0 to 12N, the fluctuation range of the $p_{\text{slip}}$ collected for each material is approximately within 2 mm. Using 2 mm as a reference, to ensure the data collected after slipping occurs is reasonable, we choose a distance of 0.9 mm (0.2 s $\times$ 4.5 mm/s) as the range.
@@
The Franka Panda robot features force-torque feedback of the end-effector, eliminating the need for additional sensors. We simultaneously record force feedback during the data collection process. Although the force feedback is not useful for model training, we can visualize its relationship with $p_{\text{slip}}$ to verify the reasonableness of the collected data, as shown in Fig. 4. The plot reveals a linear positive correlation between $p_{\text{slip}}$ and the total external force applied to the grasped object for each material, which is consistent with the model of Coulomb friction. According to the Coulomb friction model, the maximum static friction force $F_{\text{max}}=\mu F_{\text{normal}}$, where $F_{\text{normal}}$ is the normal force applied on the contact surface, and $\mu$ is the coefficient of static friction. Moreover, we can assume that $F_{\text{normal}}$ has a linear relationship with the gripper width. Thus, the plot in Fig. 4 is consistent with the aforementioned equation. This demonstrates the effectiveness of our automated data collection method.
@@
We also record some tactile raw images when not applying external force. For these images and some cases that the external force is too small, there is no slipping occurring. We run linear regression on the remaining data points of this material to estimate $p_{\text{slip}}$ for these data points, as shown in Fig. 4. We also add some tactile images without any contact to the dataset. For the images without contact, we set $p_{n}=30~{}\text{mm}$, $p_{\text{slip}}=28.5~{}\text{mm}+\text{RAND}(-0.5~{}\text{mm},0.5~{}\text{mm})$, and $v_{n}=(p_{\text{slip}}-p_{n})/3$, where $\text{RAND}(-0.5~{}\text{mm},0.5~{}\text{mm})$ generates a random distance between -0.5 mm and 0.5 mm. 3 refers to 3 seconds. This parameter controls the desired velocity of gripper width when there is no contact. A smaller value might result in an excessively high desired velocity, which may not be feasible for low-level tracking. If the value is set too high, the gripper contracts too slowly when there is no contact, resulting in a significant decrease in the overall system efficiency. Therefore, the choice of 3 seconds is also based on experiments tailored to our hardware specifications. During the training, these images without contact allow the gripper to continue closing until it makes contact with the object.
@@ III-E Loss Function
Remember that in data collection, $p_{\text{slip}}$ represents the maximum gripper width required for the gripper to hold the data collection blocks without slipping when external forces are present. Selecting $p_{\text{slip}}$ as the target position means aiming for training a model that can output the maximum gripper width that can maintain stability of the grasped object in the hand, which corresponds to applying minimal gripping force on the grasped object.
@@ III-F Controller Implementation
The model is trained using discrete data points, while actual tactile feedback is continuous and real-time. This difference in data formats poses significant challenges in deploying a trained model as a real-time controller with optimal performance. In our method, the proposed MPC layer combines both model-based and data-driven formulations, providing a way to fine-tune the model-based part for better control performance with real-time tactile feedback. To implement the trained model as a controller, we need to make three adjustments to the MPC layer, as shown in Fig. 5. Firstly, we increase the weight coefficients $Q_{v}$ and $Q_{a}$ by multiplying scalars $K_{v}$ and $K_{a}$, respectively, to increase the weights coefficients to speed up the controller’s convergence. Secondly, we incorporate an additional model-based feedback signal, the marker displacement, to enhance the response speed to the tangential contact force. Finally, we add gripper motion saturation constraints in the optimization problem to ensure that the control input and output are dynamically feasible. The implemented MPC controller is determined by solving the optimization problem
@@
The reason we add constraints during the implementation of the controller instead of incorporating them during the learning phase is that an unconstrained MPC layer can better capture the features of different objects to build more applicable dynamic models. When utilizing the learned model as a controller, the application of constraints enables the entire model to adapt better to real-time tactile feedback and generate dynamically feasible motion.
@@ IV-A Proportional Derivative Control
Firstly, we define several control signals that can be extracted from tactile images. We denote the contact area of the grasped object with the tactile sensor as $c$. The contact area $c$ can be obtained by thresholding the depth image or difference image, as illustrated in Figs. 10 and 11. Additionally, we estimate the tangential contact force by tracking the marker displacements. We represent the length of the sum of marker displacements as $d$.
It’s noteworthy that the $d$ and $c$ here are tactile signals with clear meanings, extracted by other modules, and low-dimensional. This is distinct from $\mathbf{f}_{n}$ used in LeTac-MPC. $\mathbf{f}_{n}$ is an embedding generated by the encoder, which lacks a specific explicit meaning. During end-to-end training, we anticipate that our model would learn the physical representations of grasped objects from tactile images, which is $\mathbf{f}_{n}$.

### v2

@@ III-C Training
For this MPC layer, we choose $\mathbf{A}_{f}\in\mathbb{R}^{M}$ and $\mathbf{Q}_{f}\in\mathbb{R}^{M\times M}$ as the parameters we want to learn, since they cannot be derived from the model-based formulation. On the other hand, the scalars $Q_{v}$, $Q_{a}$, and $P$ are model-based control parameters that are directly related to the response and convergence speed of the MPC. Therefore, we do not consider them as parameters to learn and can assign reasonable values directly. Additionally, it is important to note that $Q_{v}$, $Q_{a}$, and $P$ must be positive. The positivity of these scalars is crucial for the feasibility of optimization problem (III-B). We will explain why these scalars must be positive in Theorem 1.
Optimization problem (III-B) is a quadratic program (QP). The work in [43] demonstrates that a QP can be represented as a differentiable layer. In addition, a method is proposed to solve multiple QPs in batch form in [43]. Therefore, the NN shown in Fig. 1 can perform forward pass and backpropagation in batch form if the optimization problem (III-B) is feasible throughout the training process. To ensure feasibility, the resulting QP’s Hessian matrix must remain symmetric positive definite when $\mathbf{A}_{f}$ and $\mathbf{Q}_{f}$ change during training [44].
Regarding the feasibility of our proposed differentiable MPC layer, we state the following theorem.
@@
Since the optimization problem (III-B) does not have constraints besides (22), the resulting QP is unconstrained. Therefore, to prove the theorem, it is sufficient to demonstrate that the Hessian matrix of the resulting QP is symmetric positive definite [44]. The resulting Hessian matrix from (III-B) is
@@
Because $Q_{v},Q_{a},P>0$ and $\mathbf{Q}_{f}$ are symmetric positive definite, $\bar{\mathbf{Q}}$ and $\bar{\mathbf{Q}}_{a}$ are symmetric positive definite. Therefore, by equation (25), we can see that the Hessian matrix is symmetric positive semi-definite.
To prove that the Hessian matrix is symmetric positive definite, we can prove rank($\bar{\mathbf{S}}$$=N$. Because $\mathbf{B}\neq\mathbf{0}$, we can clearly see from equation (30) that rank($\bar{\mathbf{S}}$$=N$. Therefore, we have shown that the Hessian matrix of the resulting QP is symmetric positive definite, which completes the proof of the theorem. ∎
Theorem 1 demonstrates that if we ensure that $\mathbf{Q}_{f}$ is symmetric positive definite, the selection of $N$ and $M$, and uncertainty in $\mathbf{f}_{n}$ and $\mathbf{A}_{f}$ will not affect the QP’s feasibility. As a result, we can freely tune the parameters and train the model stably. To ensure that $\mathbf{Q}_{f}$ remains symmetric positive definite during training, we utilize a Cholesky factorization
@@
Although Theorem 1 shows that the selection of the prediction horizon $N$ and the embedding dimension $M$ does not affect the feasibility of the QP, it is still important to choose the appropriate values for these parameters. Very small $N$ and $M$ can negatively impact the network’s ability to extract useful information and generalize well. Conversely, setting $N$ and $M$ too large can slow down the optimization problem’s solution due to the increase in dimensionality when implementing the trained model as a controller, which can result in a low control frequency. Moreover, when $N$ is excessively large, the linear MPC layer may result in a large prediction error. Therefore, the choice of $N$ and $M$ should balance the expressiveness of the model and the computational efficiency of the controller. In our implementation, we choose $N=15$ and $M=20$ to strike this balance.
@@
Figure 2: For each type of material, bottom left is the raw tactile image overlapping with the marker tracking when grasping the corresponding block and bottom right is the raw tactile image without marker tracking. These tactile images are collected under the same grasping force. Our goal here is to show that different materials have different physical properties, which leads to different features in the tactile image under the same grasping conditions. Therefore, we do not estimate the specific values of tangential force here.
@@
[…]
Figure 3: Left: data collection setup. Right: automated data collection pipeline. The $xyz$-coordinate of the end-effector’s position increments is shown in the left figure.
@@ III-D Data Collection
We selected four different types of materials for data collection, as illustrated in Fig. 2. It is important to mention that we chose only standardized blocks without surface texture and do not require specific sizes, as long as they fit the gripper. By doing so, the data collection process is simplified. we show in Section V that our proposed network model can learn from these standardized blocks and generalize to other daily objects with different textures, sizes, shapes, and physical properties.
@@
These blocks are used as grasped objects to capture raw tactile images. The tactile feedback of these blocks varies depending on their unique physical properties. As shown in Fig. 2, we can see the differences in tactile feedback between these 4 blocks.
@@
The contact area of hard materials is smaller but more distinct, while the contact area of soft materials is larger but subtler. This is because of the difference of the objects’ stiffness. When objects with different stiffness are grasped under the same applied force, the deformation of the grasped objects and the resulting deformation on the GelSight surface will be different, resulting in different raw tactile images.
@@
The 3D printing block serves the role of a rigid object. We utilized a Markforged Onyx printer, which prints using micro carbon fiber filled nylon. This material is Onyx composite material. Its tensile modulus is 2.4 Gpa, and its flexural modulus is 3 Gpa [45]. When printed with a block form factor, it behaves rigid when grasped by our grippers. Compared with the hard block made by 3D printing, hard rubber is softer, so the contact area is larger. Additionally, the contact area of gel is the most subtle. The gel is made of XP-565 with shore A hardness of 27 [46], which is approximate 0.0009942 GPa of the Young’s modulus [47]. The stiffness of the softer rubber is between the gel and hard rubber. Finally, the difference in marker displacements can be attributed to the fact that materials with higher surface friction coefficients tend to induce a more pronounced displacement on the GelSight’s surface. It is worth noting that XP-565 is a commonly used material to make silicone gel for GelSight sensors. The gel made of XP-565 has a nominal shore A hardness of 27 [46]. Therefore, the stiffness of the gel in Fig. 2 is very close to the gel of the GelSight sensors made by XP-565. We believe that this point can serve as a benchmark for selecting data collection blocks: the chosen blocks usually possess a stiffness greater than that of the sensor gel. Otherwise, the deformation caused by the block on the sensor gel would be very subtle or even be invisible.
@@
We believe that when other people select materials to construct a dataset, following our approach of choosing one hard material (such as rigid 3D printed objects or wood) and then selecting materials with progressively lower stiffness, up to a total of four materials, would be an effective process for dataset construction.
@@
Our network input consists only of raw images. Since marker displacements are actually obtained by optical flow from raw images, the raw images themselves already contain this information. To collect data for training the NN shown in Fig. 1, we propose an automatic data collection pipeline utilizing a dual-arm setup. We use Franka Panda robotic arm. One of the arms serves as the leader arm, and we control it by making it follow specific trajectories. The other arm serves as the follower arm, and we implement an impedance controller for it. The WSG 50-110 grippers are attached to the end-effectors of both arms, with one finger on the follower arm mounting the GelSight tactile sensor.
@@
Next, the follower gripper gradually relaxes with a velocity of 4.5 mm/s, and we record the width of the follower gripper $p_{n}$ and the raw image of the tactile feedback at 60 Hz. When the follower gripper’s grasping force applied to the object cannot provide a maximum static friction force greater than the external force from the impedance controller, slipping occurs between the follower gripper and the object. This results in the follower arm moving back to the equilibrium point of the impedance controller. We record the width of the follower gripper when slipping occurs as the label for all raw images recorded during the entire trial, denoted as $p_{\text{slip}}$.
Ideally, we hope that the leader gripper maintains an absolutely tight grasping throughout the process and there is no slipping between the leader gripper and the grasped object. To achieve that, we maximize the grasping force of the leader gripper on the data collection block and enhance the friction between the leader gripper’s fingers and the block by adding a gel layer on the finger. Naturally, we could not guarantee that there would be no movement between the object and the leader gripper during extended periods of data collection. However, our setup ensures that any movement, if it occurred, is minimal and does not affect data collection. Therefore, when the leader gripper stays static, the grasped object stays static and we set the velocity of the end-effector of the follower arm to be greater than 5 mm/s as the condition for indicating slipping. Once slipping occurs for 0.2 s, we stop recording the raw images. The reasons of selecting 0.2 s are as follows. For each material, as shown in Fig. 4, when the external force varies from 0 to 12N, the fluctuation range of the $p_{\text{slip}}$ collected for each material is approximately within 2 mm. Using 2 mm as a reference, to ensure the data collected in a reasonable range (for example, to avoid collecting too many images without contact), we chose a distance of 0.9 mm (0.2 s $\times$ 4.5 mm/s) as the gripper width range of one data collection trial. It is worth noting that the choice of 0.2 s here is approximate. The goal is to prevent the collected data from becoming overly biased.
@@
The Franka Panda robot features force-torque feedback of the end-effector, eliminating the need for additional sensors. We simultaneously record force feedback during the data collection process. Although force feedback is not useful for model training, we can visualize its relationship with $p_{\text{slip}}$ to verify the reasonableness of the collected data, as shown in Fig. 4. Fig. 4 shows that the gripper width and the external force applied to the block are approximately negatively correlated. This indicates an approximate linear positive correlation between the pressure and the total external force applied to each block, aligned with the Coulomb friction model. According to this model, the maximum static friction force $F_{\text{max}}=\mu F_{\text{normal}}$, where $F_{\text{normal}}$ is the normal force exerted on the contact surface and $\mu$ is the coefficient of static friction. Additionally, we know that the gripper width and $F_{\text{normal}}$ are negatively correlated (a smaller gripper width means tighter grasping and a larger $F_{\text{normal}}$). Therefore, the depiction in Fig. 4 corroborates this equation. This demonstrates the effectiveness of our automated data collection method.
@@
We also record some raw tactile images when not applying external force. For these images and some cases where the external force is too small, there is no slipping occurring. We run a linear regression on the remaining data points of this material to estimate $p_{\text{slip}}$ for these data points, as shown in Fig. 4. We also add some tactile images without any contact to the dataset. For images without contact, we set $p_{n}=30~{}\text{mm}$, $p_{\text{slip}}=28.5~{}\text{mm}+\text{RAND}(-0.5~{}\text{mm},0.5~{}\text{mm})$, and $v_{n}=(p_{\text{slip}}-p_{n})/3$, where $\text{RAND}(-0.5~{}\text{mm},0.5~{}\text{mm})$ generates a random distance between -0.5 mm and 0.5 mm. 3 refers to 3 seconds. This parameter controls the desired velocity of gripper width when there is no contact. A lower value might result in an excessively high desired velocity, which may not be feasible for low-level tracking. If the value is set too high, the gripper contracts too slowly when there is no contact, resulting in a significant decrease in overall system efficiency. Therefore, the choice of 3 seconds is also based on experiments tailored to our hardware specifications. During training, these images without contact allow the gripper to continue closing until it makes contact with the object.
We mount only one GelSight sensor on the follower gripper, and the other finger is with the same shape and no sensor mounted. Similarly, the two fingers on the leader gripper are all dummy fingers without mounted sensors since the leader arm is used to provide external forces during data collection. The reason for using only one GelSight on one side is because a single GelSight sensor suffices to furnish LeTac-MPC with adequate information for effective grasping. For the parallel gripper, the GelSight sensors mounted on different fingers can offer detailed information on the shape and texture of an grasped object’s opposing sides, but often provide overlapping information about the information of contact force, and states and properties of the grasped object. GelSight only can perceive a very small local contact area of the object, and the idea of LeTac-MPC is using this small local area (raw tactile image) to estimate the grasping state and object property implicitly and then generate the corresponding actions. Therefore, we only mount one GelSight sensor, both for data collection and controller deployment. In addition, adding a redundant GelSight sensor would increase the computational load for real-time processing, subsequently reducing the control frequency. In addition, we would like to highlight the following papers [16, 48]. Both of them have only one finger equipped with GelSight sensor but can accomplish complex manipulation tasks.
@@ III-E Loss Function
Recall that in data collection, $p_{\text{slip}}$ represents the maximum gripper width required for the gripper to hold the data collection blocks without slipping when external forces are present. Selecting $p_{\text{slip}}$ as the target position means aiming for training a model that can output the maximum gripper width that can maintain stability of the grasped object in the hand, which corresponds to applying minimal gripping force to the grasped object.
@@ III-F Controller Implementation
The model is trained using discrete data points, while the actual tactile feedback is continuous and real-time. This difference in data formats poses significant challenges in deploying a trained model as a real-time controller with optimal performance. In our method, the proposed MPC layer combines both model-based and data-driven formulations, providing a way to fine-tune the model-based part for better control performance with real-time tactile feedback. To implement the trained model as a controller, we need to make three adjustments to the MPC layer, as shown in Fig. 5. Firstly, we increase the weight coefficients $Q_{v}$ and $Q_{a}$ by multiplying the scalars $K_{v}$ and $K_{a}$, respectively, to increase the weights coefficients to speed up the controller’s convergence. Secondly, we incorporate an additional model-based feedback signal, the marker displacement, to enhance the response speed to the tangential contact force. Finally, we add gripper motion saturation constraints in the optimization problem to ensure that control input and output are dynamically feasible. The implemented MPC controller is determined by solving the optimization problem
@@
The reason we add constraints during the implementation of the controller instead of incorporating them during the learning phase is that an unconstrained MPC layer can better capture the features of different objects to build more applicable dynamic models. When utilizing the learned model as a controller, the application of constraints enables the entire model to better adapt to real-time tactile feedback and generate dynamically feasible motion.
@@ IV-A Proportional Derivative Control
Firstly, we define several control signals that can be extracted from tactile images. We denote the contact area of the grasped object with the tactile sensor as $c$. The contact area $c$ can be obtained by thresholding the depth image or the difference image, as illustrated in Figs. 10 and 11. Additionally, we estimate the tangential contact force by tracking the marker displacements. We represent the length of the sum of marker displacements as $d$.
It should be noted that $d$ and $c$ here are tactile signals with clear meanings, extracted by other modules, and low-dimensional. This is distinct from $\mathbf{f}_{n}$ used in LeTac-MPC. $\mathbf{f}_{n}$ is an embedding generated by the encoder, which lacks a specific explicit meaning. During end-to-end training, we anticipate that our model would learn the physical representations of grasped objects from tactile images, which is $\mathbf{f}_{n}$.